OpenStudy (anonymous):

The number of real roots of the equation |x^2+x-6| + 2|x| - 4 =0 is?

OpenStudy (anonymous):

@satellite73 @Hero @hartnn

OpenStudy (anonymous):

I'm guessing six right now. I'll have to play with it a little more to be sure.

OpenStudy (anonymous):

U see @CliffSedge " REAL ROOTS"

OpenStudy (anonymous):

I know, I know, that's just a first approximation to set some bounds . . .

OpenStudy (anonymous):

The Options are 0 1 2 3

OpenStudy (anonymous):

Hmm, must have done something wrong, because now I am getting 8 solutions. I'll have to check them with the original equation to see if some are erroneous.

OpenStudy (raden):

if i use trial error, i get 2

hero (hero):

Hint: Use the rule \[|a| =\sqrt{ a^2}\] to solve

OpenStudy (anonymous):

I know x=2 works by inspection. I only found one other that works. I'll also say only 2 solutions.

OpenStudy (anonymous):

The answer shuld be 1 solution.....)

OpenStudy (anonymous):

\[|a| = \pm a\]

OpenStudy (anonymous):

|x^2+x-6| + 2|x| - 4 =0 x^2 + x - 6 + 2x - 4 =0 x^3 + 3x - 10 =0 --------- 1 |x^2+x-6| + 2|x| - 4 =0 -x^2-x+6 - 2x - 4 =0 x^2 + 3x - 2 =0 ----------2

OpenStudy (raden):

for x>0, 1st equation should be : x^2 + 3x - 10 = 0 (x+5)(x-2) = 0 x=2 is which satisfy be a solution

OpenStudy (anonymous):

I see where I made an error (lost track of a negative sign . . .). Now I only see x=2 as solution.

hartnn (hartnn):

if u substitute back the other 3 roots u got in original equation, it will not get satisfied. so x=2 is the only real solution